OpenStudy (anonymous):

The value of a rare painting has increased each year since it was found at a garage sale. The value of the painting is modeled by the function f(x) = 799(1.03)x. What does the 799 represent? What will the painting be worth after 5 years? Round your answer to the nearest dollar.

OpenStudy (anonymous):

@CBARREDO1 @campbell_st @MathGuyThatWantsHelp

OpenStudy (cbarredo1):

799 represents the value of the painting when it was found; the painting will be worth $926 after 5 years

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (cbarredo1):

Pretty sure

OpenStudy (campbell_st):

well one of the key principals to this site is to help understanding and not give answers...

OpenStudy (campbell_st):

if x is time in years.... this growth model is just an application of the compound interest formula. so if you subsitute x = 0 that with give the initial value of the painting... remember anything to the power zero is 1

OpenStudy (campbell_st):

the 2nd part of the question tells you 5 years later so x =5 make the substitution and calculate the value... \[f(5) = 799 \times (1.03)^5\] that's all that is needed

OpenStudy (anonymous):

can u help me with one more @CBARREDO1

OpenStudy (anonymous):

For f(x) = 0.01(2)x, find the average rate of change from x = 12 to x = 15.

OpenStudy (cbarredo1):

Sure

OpenStudy (anonymous):

For f(x) = 0.01(2)x, find the average rate of change from x = 12 to x = 15.